On July 1, 2016, Killearn Company acquired 96,000 of the outstanding shares of Shaun Company for $17 per share. This acquisition gave Killearn a 40 percent ownership of Shaun and allowed Killearn to significantly influence the investee's decisions.

As of July 1, 2016, the investee had assets with a book value of $4 million and liabilities of $1,063,500. At the time, Shaun held equipment appraised at $262,500 above book value; it was considered to have a seven-year remaining life with no salvage value. Shaun also held a copyright with a five-year remaining life on its books that was undervalued by $707,500. Any remaining excess cost was attributable to goodwill. Depreciation and amortization are computed using the straight-line method. Killearn applies the equity method for its investment in Shaun.

Shaun's policy is to declare and pay a $1 per share cash dividend every April 1 and October 1. Shaun's income, earned evenly throughout each year, was $588,000 in 2016, $620,800 in 2017, and $667,400 in 2018.

In addition, Killearn sold inventory costing $99,600 to Shaun for $166,000 during 2017. Shaun resold $85,000 of this inventory during 2017 and the remaining $81,000 during 2018.

  1. Determine the equity income to be recognized by Killearn during each of these years.

  2. Compute Killearn's investment in Shaun Company's balance as of December 31, 2018.

A) Equity income 2016 -

Equity income 2017 -

Equity income 2018 -

B) Investment in Shaun
Answer #1

Please find below table useful to compute desired results: -

H Equity Income 2016 Basic Equity Accrual =B12*B3*1/2 Amortization =D28*1/2 Equity Income 2016 =E2-E3 Equity Income 2017 Basi

End results would be as follows: -

F 96,000 17 40% 4,000,000 1,063,500 262,500 Equity Income 2016 Basic Equity Accrual Amortization Equity Income 2016 117,600 3